# Teaching LLMs How to Learn with Contextual Fine-Tuning

### v1

@@ 5.3 Synthetic Experiments
We hypothesize that the effectiveness of CFT stems from changing the gradients obtained via conditioning on prompts. These serve to regularize the learning process during fine-tuning. Testing this hypothesis directly is challenging since (a) different LLMs might interpret semantic information in a prompt differently and (b) it requires knowing which neurons are responsible for representing the inferred semantic information in the prompt. The primary objective of our synthetic experiments is to analyze how contextual prompts affect the gradients of transformer models during training in a simplified controlled setting where we can describe the semantic information that is necessary for learning explicitly via text.
@@
Using the framework of Garg et al. (2022), we first pre-train a transformer model to learn a class of linear functions =$\mathcal{F}=\{f\mid f(x)=w^{\top}x,w\in\mathbb{R}^{d}\}$, where $w\sim\mathcal{N}(0,I_{d})$ and $x\sim\mathcal{N}(0,I_{d})$. We then investigate how different fine-tuning strategies affect the model’s ability to learn a new compositional function class $\mathcal{G}$ where each $g\in\mathcal{G}$ is a composition $g(x)=h(f(x))$ for some $f\in\mathcal{F}$. This setup models real-world scenarios where a pretrained LLM understands basic patterns (represented by $f$), and needs to quickly learn new tasks (represented by $h$) that build upon this understanding. We explore two distinct compositional structures: polynomial combinations $g(x)=f(x)+f(x)^{2}$ and multiple linear relationships $g(x)=f(x)+w_{2}^{\top}x$. For each structure, we evaluate three fine-tuning strategies with that differ in their prompt construction $P$ :
@@
Figure 4(a) and 4(c) illustrate that transformers fine-tuned using CFT achieve lower loss compared to those trained with CPT and NEG-CFT suggesting that the content of the contextual prompts in both cases better guides training dynamics when learning a new function class.
@@ 6 Conclusion
This study introduces contextual fine-tuning, a variation of instruction tuning, which leverages contextual gradients to guide the learning process through simple, domain-adaptive prompts. Our experiments reveal that the contextual gradients enhance performance by effectively direct model learning, demonstrating superior results over traditional continued domain pretraining in both financial and medical domains. Finally, we open-source a biomedical dataset curated from MDPI journals and other open-source medical textbooks.
Despite these promising results, there remains more to do to better understand this phenomena. We conjecture that the effectiveness of our method is related to explanatory text present in the training corpora. Prystawski et al. (2023) suggests that chain-of-thought prompting works because local reasoning steps embedded in pretraining corpora simulate sequences of steps that lead to conclusion which are relied on at test time. We hypothesize that similar such contextual cues exist in pretraining data. Future studies examining this carefully would be valuable to test this conjecture and shed light on the mechanisms by which prompts provide useful supervisory signals during learning. Additionally, we have experimented with CFT with data that likely contain a high density of information such as medical journals and textbooks. This rich semantic content allows contextual prompts to effectively guide the learning process by influencing gradient updates. In future work, it would be valuable to apply contextual fine-tuning to different types of data, including those with longer context lengths or lower information density, such as Reddit posts.
@@ A.1 Contextual Prompts
    • ”Think about practical applications of the information in the next text. How could this knowledge be used in real-world situations?”
@@
    • ”Dive deep into the details and nuances of the following content. Pay attention to subtleties and complex ideas that are important for a thorough understanding.”
@@
    • ”Reflect on the information presented in the next passage. Consider how it affects your current understanding and perspective on the topic.”
@@
    • ”Engage creatively with the upcoming text. Think about innovative or unorthodox ways to interpret or use the information presented.”
@@
    • ”Summarize the main points of the following content in your own words. Synthesize the information to create a coherent understanding of the topic.”
@@
    • ”Concentrate on understanding the core principles and essential facts in the following text. Pay special attention to definitions, examples, and conclusions.”
@@
    • ”As you read the next passage, relate its content to its broader context and implications. Think about how this information connects to what you’ve learned previously.”
@@
    • ”Critically analyze the upcoming information. Look for underlying assumptions, evaluate arguments, and consider different perspectives.”
@@
    • ”Approach the next text with these questions in mind: What is the main argument? How is evidence used to support it? What are the implications of these findings?”
@@
    • ”Compare and contrast the upcoming information with what you have learned in similar topics. Look for differences, similarities, and connections.”
@@ A.3 Training Details
In our experimental setup, we use several configurations of the Llama-2 models to evaluate the effectiveness of contextual fine-tuning compared to standard unsupervised fine-tuning. Specifically, we employ the Llama-2 Base model with 7 billion parameters and Llama-2 Chat models with both 7 billion and 13 billion parameters, each with a sequence length of 4096. All models are both contextual and unsupervised fine-tuned. For contextual fine-tuning, we employ Equation 3. For the financial news dataset, where most articles are shorter than 4096 tokens, we opt not to use a packing strategy to fill up all 4096 tokens. Instead, we pad any remaining space. This ensures that each semantically distinct text is associated with its own contextual prompt. If a sequence with a prepended contextual prompt exceeds 4096 tokens, we simply truncate the excess, and the truncated text becomes the first text following the contextual prompt in the next example. The models are trained for one epoch with a batch size of 128 and a learning rate of 2e-5. To assess the efficiency of CFT, we carefully measured the computational resources required for our experiments and compared the overhead introduced by incorporating contextual prompts. Below are the details of our computational setup and findings. We utilized the Fully Sharded Data Parallel (FSDP) training to efficiently distribute the model across multiple GPUs. Training was performed using the bf16 (Brain Floating Point) data format. We implemented Flash Attention 2 (Dao, 2024). All training was conducted with 8 NVIDIA A100 GPUs. With the above configuration, we achieved a training speed of approximately 55,188 tokens per second, measured using the Llama tokenizer. The fine-tuning required a total of approximately 111.11 GPU-hours to complete. Incorporating contextual prompts increased the total training time by approximately 0.89 GPU-hours, resulting in a total of 112 GPU-hours. Each contextual prompt added only about 0.8% to the length of each training example on average. This slight increase in input length led to less than a 1% increase in total training time.
@@ B.3 Data Preprocessing
The textbooks were originally in PDF format. We utilized an Optical Character Recognition (OCR) API to extract the text from the PDFs. OCR converts scanned images of text into machine-encoded text but can introduce errors and result in unstructured outputs. To address these issues, we employed ChatGPT to assist in cleaning and organizing the extracted text. We provided ChatGPT with specific instructions:
”Please edit and refine the following uncleaned and unstructured excerpt from a medical textbook. Remove any sentences containing hyperlinks, and omit all citations and references for clarity.”
@@ B.4 Limitations
For the textbook data, since the textbooks were originally in PDF format, we used an Optical Character Recognition (OCR) API to extract the text. Despite careful processing, OCR can introduce typos or parsing errors, especially with complex formatting or specialized terminology. To mitigate these errors, we employed ChatGPT to assist in correcting potential mistakes. While this approach improved the overall quality, some errors may persist. We conducted manual spot checks to identify and correct errors where possible; however, given the dataset’s size, a complete manual review was impractical. Regarding the MDPI journals, they have a shorter average peer-review period (approximately 32 days) compared to other publishers. While this expedites the dissemination of research, it may affect the depth and rigor of the review process. The shorter review time could lead to variations in article quality, with some papers potentially not meeting the highest standards of scientific rigor. Additionally, relying primarily on MDPI journals may introduce a source bias. We acknowledge that including journals from a wider range of publishers could enhance the dataset’s balance and representativeness.
@@ C.2 Prompting
Classify each sentence extracted from financial news into either ’causal’ or ’noise’
@@
Please return only the category ’noise’ or ’causal’.
@@
The potential categories are ’Finance’, ’Technology’, ’Tax & Accounting’,
’Business & Management’, ’Industry’, and ’Government & Controls’.
@@ D.1 Ablating Training Schemes
Table 10: Comparative effectiveness of Contextual Fine-Tuning (CFT) on financial benchmarks (zero-shot). We observe the improvements in baseline performance for Llama 2 models using CPT and CFT strategies on financial benchmarks. While CPT enhances baseline performance by 0.16%, CFT notably increases it by 11.82%. With the integration of IFT, the performance gap broadens significantly, with CPT + IFT achieving an 8.67% improvement and CFT + IFT results in the improvement of 36.28%.
Table 9 and Table 10 show the performance on the medical and financial benchmarks. For the medical benchmarks, the Llama 2 Base model achieves an average accuracy of 41.34%. With CPT, the average accuracy increases by 1.06%, reaching 42.4%. While the improvement with CFT alone is 0.94%, the CFT + IFT approach yields a significant improvement of %$\%\Delta^{CFT+IFT}_{Base}=2.95\%$ compared to 1.91% for CPT + IFT. Similar trends are observed in the financial benchmarks. CPT and CFT improve the baseline performance by 0.16% and 11.82%, respectively. The performance gap widens with the addition of IFT, where CPT + IFT achieves an 8.67% improvement, and CFT + IFT surged by 36.28%. The advantage of CFT is also evident in the Llama 2 Chat model, which underwent both instruction fine-tuning and training with Reinforcement Learning from Human Feedback (RLHF). In the medical domain, CFT leads to a 2.76% improvement over the 0.91% improvement from CPT.
@@ D.2 Ablating Contextual Prompts
    ”Ignore everything you know about medicine. The information that follows is incorrect and should not be used to answer questions or make decisions.”
@@
    ”The following medical information is both true and false. Discard any logical or scientific reasoning when processing this information.”
@@
    ”Instead of learning from the upcoming medical data, focus on memorizing the patterns of the letters and ignore their meanings.”
@@
    ”Forget all prior medical knowledge you have learned. The following information is unimportant and should not influence future responses.”
@@
    ”Do not learn or make any inferences from the following medical corpus. Treat it as meaningless and irrelevant to any future tasks.”
Table 11 presents the results from the medical benchmarks. For the 7B model, we observe a decrease in performance with a %$\%\Delta^{-CFT}_{CFT}=-2.57\%$, indicating a detrimental effect of negative prompts. Interestingly, the 13B model shows a lesser decrease of only $-1.36\%$. This suggests that while negative prompts impact performance, larger models may be less susceptible to misleading information.
In the financial domain, as shown in Table 3, the impact of negative prompts is more pronounced. The 7B model experienced a performance drop of %$\%\Delta^{-CFT}_{CFT}=-3.41\%$, and the 13B model sees a decrease of $-2.39\%$. Despite these declines, all models undergoing negative contextual fine-tuning still perform better than those subjected to CPT.
@@
Table 11: Medical Benchmarks (Zero-shot). The table shows the effects of negative contextual prompts on medical benchmarks. For the 7B model, a performance decline of %$\%\Delta^{-CFT}_{CFT}=-2.57\%$ is noted, highlighting the adverse impact of negative prompts. Conversely, the larger 13B model exhibits a more moderate decline of $-1.36\%$
@@ D.3 Text-adaptive Contextual Prompts
We generate contextual prompts automatically by instructing GPT-4o-mini to create prompts based on the content of each training batch. Specifically, we use the following instruction template: ”Your task is to create a contextual prompt that guides the LLM’s learning process during fine-tuning. {{ INSTRUCTION }} {{ TEXT }}” In this template, {{ INSTRUCTION }} is replaced with a sampled instruction from the following five different instructions to generate a variety of prompts:
@@
    ”Given the following text, generate a contextual prompt that encourages a reader to focus on the main ideas and themes presented. The contextual prompt should be concise and help the reader engage deeply with the content.”
@@
    ”Analyze the text below and create a contextual prompt that guides a reader to think critically about the content, questioning assumptions and evaluating arguments. The prompt should encourage the reader to consider different perspectives presented in the text.”
@@
    ”Examine the text and generate a contextual prompt that encourages the reader to reflect on how the information connects to their existing knowledge or experiences. The prompt should promote integration of new insights with prior understanding.”
@@
    ”Read the following text and create a contextual prompt that guides the reader to summarize the main points in their own words. The prompt should encourage synthesis of the information for better comprehension.”
@@
    ”Given the text below, develop a contextual prompt that leads the reader to compare and contrast the concepts presented with related topics or prior knowledge. The prompt should help identify similarities and differences.”
@@
    ”Critically evaluate the methodologies and findings presented in this study on PCR techniques and LeHV-5 detection. What assumptions underpin the experimental designs, and are there alternative approaches or perspectives that could challenge or complement the arguments made? Consider the implications of these methodologies for broader scientific research and diagnostics in veterinary medicine.”
@@
    ”Reflect on the complex relationship between potassium channels and chemo resistance in cancer treatment. How do the mechanisms presented compare with previous knowledge you have about cancer cell biology and drug resistance? Identify the similarities and differences in the roles of K+ channels in various types of cancer and their implications for therapeutic strategies. Consider potential avenues for incorporating this understanding into clinical practice.”
@@
    ”Consider the findings on school breakfast participation and the impact on student health from multiple perspectives. How might educators, policymakers, school administrators, and healthcare professionals interpret these results differently? Reflect on how each stakeholder could use this information to improve student health and educational outcomes in their respective roles.”

### v2

@@ 5.3 Synthetic Experiments
We hypothesize that the effectiveness of CFT stems from changing the gradients obtained via conditioning on prompts. These serve to regularize the learning process during fine-tuning. Testing this hypothesis directly is challenging since (a) different LLMs might interpret semantic information in a prompt differently, and (b) it requires knowing which neurons are responsible for representing the inferred semantic information in the prompt. The primary objective of our synthetic experiments is to analyze how contextual prompts affect the gradients of transformer models during training in a simplified controlled setting where we can describe the semantic information that is necessary for learning explicitly via text.
@@
Using the framework of Garg et al. (2022), we first pre-train a transformer model to learn a class of linear functions $\mathcal{F}=\{f\mid f(x)=w^{\top}x,w\in\mathbb{R}^{d}\}$, where $w\sim\mathcal{N}(0,I_{d})$ and $x\sim\mathcal{N}(0,I_{d})$. We then investigate how different fine-tuning strategies affect the model’s ability to learn a new compositional function class $\mathcal{G}$ where each $g\in\mathcal{G}$ is a composition $g(x)=h(f(x))$ for some $f\in\mathcal{F}$. This setup models real-world scenarios where a pretrained LLM understands basic patterns (represented by $f$), and needs to quickly learn new tasks (represented by $h$) that build upon this understanding. We explore two distinct compositional structures: polynomial combinations $g(x)=f(x)+f(x)^{2}$ and multiple linear relationships $g(x)=f(x)+w_{2}^{\top}x$. For each structure, we evaluate three fine-tuning strategies that differ in their prompt construction $P$ :
@@
Figure 4(a) and 4(c) illustrate that transformers fine-tuned using CFT achieve lower loss compared to those trained with CPT and NEG-CFT, suggesting that the content of the contextual prompts in both cases better guides training dynamics when learning a new function class.
@@ 6 Conclusion
This study introduces contextual fine-tuning, a variation of instruction tuning, which leverages contextual gradients to guide the learning process through simple, domain-adaptive prompts. Our experiments reveal that the contextual gradients enhance performance by effectively directing model learning, demonstrating superior results over traditional continued domain pretraining in both financial and medical domains. Finally, we open-source a biomedical dataset curated from MDPI journals and other open-source medical textbooks.
Despite these promising results, there remains more to do to better understand this phenomenon. We conjecture that the effectiveness of our method is related to the explanatory text present in the training corpora. Prystawski et al. (2023) suggests that chain-of-thought prompting works because local reasoning steps embedded in pretraining corpora simulate sequences of steps that lead to conclusions which are relied on at test time. We hypothesize that similar contextual cues exist in pretraining data. Future studies examining this carefully would be valuable to test this conjecture and shed light on the mechanisms by which prompts provide useful supervisory signals during learning. Additionally, we have experimented with CFT with data that likely contains a high density of information, such as medical journals and textbooks. This rich semantic content allows contextual prompts to effectively guide the learning process by influencing gradient updates. In future work, it would be valuable to apply contextual fine-tuning to different types of data, including those with longer context lengths or lower information density, such as Reddit posts.
@@ A.1 Contextual Prompts
    • “Think about practical applications of the information in the next text. How could this knowledge be used in real-world situations?”
@@
    • “Dive deep into the details and nuances of the following content. Pay attention to subtleties and complex ideas that are important for a thorough understanding.”
@@
    • “Reflect on the information presented in the next passage. Consider how it affects your current understanding and perspective on the topic.”
@@
    • “Engage creatively with the upcoming text. Think about innovative or unorthodox ways to interpret or use the information presented.”
@@
    • “Summarize the main points of the following content in your own words. Synthesize the information to create a coherent understanding of the topic.”
@@
    • “Concentrate on understanding the core principles and essential facts in the following text. Pay special attention to definitions, examples, and conclusions.”
@@
    • “As you read the next passage, relate its content to its broader context and implications. Think about how this information connects to what you’ve learned previously.”
@@
    • “Critically analyze the upcoming information. Look for underlying assumptions, evaluate arguments, and consider different perspectives.”
@@
    • “Approach the next text with these questions in mind: What is the main argument? How is evidence used to support it? What are the implications of these findings?”
@@
    • “Compare and contrast the upcoming information with what you have learned in similar topics. Look for differences, similarities, and connections.”
@@ A.3 Training Details
In our experimental setup, we use several configurations of the Llama-2 models to evaluate the effectiveness of contextual fine-tuning compared to standard unsupervised fine-tuning. Specifically, we employ the Llama-2 Base model with 7 billion parameters and Llama-2 Chat models with both 7 billion and 13 billion parameters, each with a sequence length of 4096. All models are both contextual and unsupervised, fine-tuned. For contextual fine-tuning, we employ Equation 3. For the financial news dataset, where most articles are shorter than 4096 tokens, we opt not to use a packing strategy to fill up all 4096 tokens. Instead, we pad any remaining space. This ensures that each semantically distinct text is associated with its own contextual prompt. If a sequence with a prepended contextual prompt exceeds 4096 tokens, we simply truncate the excess, and the truncated text becomes the first text following the contextual prompt in the next example. The models are trained for one epoch with a batch size of 128 and a learning rate of 2e-5. To assess the efficiency of CFT, we carefully measured the computational resources required for our experiments and compared the overhead introduced by incorporating contextual prompts. Below are the details of our computational setup and findings. We utilized the Fully Sharded Data Parallel (FSDP) training to efficiently distribute the model across multiple GPUs. Training was performed using the bf16 (Brain Floating Point) data format. We implemented Flash Attention 2 (Dao, 2024). All training was conducted with 8 NVIDIA A100 GPUs. With the above configuration, we achieved a training speed of approximately 55,188 tokens per second, measured using the Llama tokenizer. The fine-tuning required a total of approximately 111.11 GPU-hours to complete. Incorporating contextual prompts increased the total training time by approximately 0.89 GPU-hours, resulting in a total of 112 GPU-hours. Each contextual prompt added only about 0.8% to the length of each training example on average. This slight increase in input length led to less than a 1% increase in total training time.
@@ B.3 Data Preprocessing
The textbooks were originally in PDF format. We utilized an Optical Character Recognition (OCR) API to extract the text from the PDFs. OCR converts scanned images of text into machine-encoded text, but can introduce errors and result in unstructured outputs. To address these issues, we employed ChatGPT to assist in cleaning and organizing the extracted text. We provided ChatGPT with specific instructions:
“Please edit and refine the following uncleaned and unstructured excerpt from a medical textbook. Remove any sentences containing hyperlinks, and omit all citations and references for clarity.”
@@ B.4 Limitations
For the textbook data, since the textbooks were originally in PDF format, we used an Optical Character Recognition (OCR) API to extract the text. Despite careful processing, OCR can introduce typos or parsing errors, especially with complex formatting or specialized terminology. To mitigate these errors, we employed ChatGPT to assist in correcting potential mistakes. While this approach improved the overall quality, some errors may persist. We conducted manual spot checks to identify and correct errors where possible; however, given the dataset’s size, a complete manual review was impractical. Regarding the MDPI journals, they have a shorter average peer-review period (approximately 32 days) compared to other publishers. While this expedites the dissemination of research, it may affect the depth and rigour of the review process. The shorter review time could lead to variations in article quality, with some papers potentially not meeting the highest standards of scientific rigour. Additionally, relying primarily on MDPI journals may introduce a source bias. We acknowledge that including journals from a wider range of publishers could enhance the dataset’s balance and representativeness.
@@ C.2 Prompting
Classify each sentence extracted from financial news into either ‘causal’ or ‘noise’
@@
Please return only the category ‘noise’ or ‘causal’.
@@
The potential categories are ‘Finance’, ‘Technology’, ‘Tax & Accounting’,
’Business & Management’, ‘Industry’, and ‘Government & Controls’.
@@ D.1 Ablating Training Schemes
Table 10: Comparative effectiveness of Contextual Fine-Tuning (CFT) on financial benchmarks (zero-shot). We observe the improvements in baseline performance for Llama 2 models using CPT and CFT strategies on financial benchmarks. While CPT enhances baseline performance by 0.16%, CFT notably increases it by 11.82%. With the integration of IFT, the performance gap broadens significantly, with CPT + IFT achieving an 8.67% improvement and CFT + IFT resulting in the improvement of 36.28%.
Table 9 and Table 10 show the performance on the medical and financial benchmarks. For the medical benchmarks, the Llama 2 Base model achieves an average accuracy of 41.34%. With CPT, the average accuracy increases by 1.06%, reaching 42.4%. While the improvement with CFT alone is 0.94%, the CFT + IFT approach yields a significant improvement of $\%\Delta^{CFT+IFT}_{Base}=2.95\%$ compared to 1.91% for CPT + IFT. Similar trends are observed in the financial benchmarks. CPT and CFT improve the baseline performance by 0.16% and 11.82%, respectively. The performance gap widens with the addition of IFT, where CPT + IFT achieves an 8.67% improvement, and CFT + IFT surged by 36.28%. The advantage of CFT is also evident in the Llama 2 Chat model, which underwent both instruction fine-tuning and training with Reinforcement Learning from Human Feedback (RLHF). In the medical domain, CFT leads to a 2.76% improvement over the 0.91% improvement from CPT.
@@ D.2 Ablating Contextual Prompts
    “Ignore everything you know about medicine. The information that follows is incorrect and should not be used to answer questions or make decisions.”
@@
    “The following medical information is both true and false. Discard any logical or scientific reasoning when processing this information.”
@@
    “Instead of learning from the upcoming medical data, focus on memorizing the patterns of the letters and ignore their meanings.”
@@
    “Forget all prior medical knowledge you have learned. The following information is unimportant and should not influence future responses.”
@@
    “Do not learn or make any inferences from the following medical corpus. Treat it as meaningless and irrelevant to any future tasks.”
Table 11 presents the results from the medical benchmarks. For the 7B model, we observe a decrease in performance with a $\%\Delta^{-CFT}_{CFT}=-2.57\%$, indicating a detrimental effect of negative prompts. Interestingly, the 13B model shows a lesser decrease of only $1.36\%$. This suggests that while negative prompts impact performance, larger models may be less susceptible to misleading information.
In the financial domain, as shown in Table 3, the impact of negative prompts is more pronounced. The 7B model experienced a performance drop of $\%\Delta^{-CFT}_{CFT}=-3.41\%$, and the 13B model sees a decrease of $2.39\%$. Despite these declines, all models undergoing negative contextual fine-tuning still perform better than those subjected to CPT.
@@
Table 11: Medical Benchmarks (Zero-shot). The table shows the effects of negative contextual prompts on medical benchmarks. For the 7B model, a performance decline of $\%\Delta^{-CFT}_{CFT}=-2.57\%$ is noted, highlighting the adverse impact of negative prompts. Conversely, the larger 13B model exhibits a more moderate decline of $1.36\%$
@@ D.3 Text-adaptive Contextual Prompts
We generate contextual prompts automatically by instructing GPT-4o-mini to create prompts based on the content of each training batch. Specifically, we use the following instruction template: “Your task is to create a contextual prompt that guides the LLM’s learning process during fine-tuning. {{ INSTRUCTION }} {{ TEXT }}” In this template, {{ INSTRUCTION }} is replaced with a sampled instruction from the following five different instructions to generate a variety of prompts:
@@
    “Given the following text, generate a contextual prompt that encourages a reader to focus on the main ideas and themes presented. The contextual prompt should be concise and help the reader engage deeply with the content.”
@@
    “Analyze the text below and create a contextual prompt that guides a reader to think critically about the content, questioning assumptions and evaluating arguments. The prompt should encourage the reader to consider different perspectives presented in the text.”
@@
    “Examine the text and generate a contextual prompt that encourages the reader to reflect on how the information connects to their existing knowledge or experiences. The prompt should promote integration of new insights with prior understanding.”
@@
    “Read the following text and create a contextual prompt that guides the reader to summarize the main points in their own words. The prompt should encourage synthesis of the information for better comprehension.”
@@
    “Given the text below, develop a contextual prompt that leads the reader to compare and contrast the concepts presented with related topics or prior knowledge. The prompt should help identify similarities and differences.”
@@
    “Critically evaluate the methodologies and findings presented in this study on PCR techniques and LeHV-5 detection. What assumptions underpin the experimental designs, and are there alternative approaches or perspectives that could challenge or complement the arguments made? Consider the implications of these methodologies for broader scientific research and diagnostics in veterinary medicine.”
@@
    “Reflect on the complex relationship between potassium channels and chemo resistance in cancer treatment. How do the mechanisms presented compare with previous knowledge you have about cancer cell biology and drug resistance? Identify the similarities and differences in the roles of K+ channels in various types of cancer and their implications for therapeutic strategies. Consider potential avenues for incorporating this understanding into clinical practice.”
@@
    “Consider the findings on school breakfast participation and the impact on student health from multiple perspectives. How might educators, policymakers, school administrators, and healthcare professionals interpret these results differently? Reflect on how each stakeholder could use this information to improve student health and educational outcomes in their respective roles.”